OpenStudy (anonymous):

C4H5O5 + 3NaHCO3 -> C4H2Na3O5+ 3H2O + 3CO2 Information: n(NaHCO3)= 5,94 mol Maleic acid (MW: 134 g/mol) used 1.1 times more of the substance quantity of the acid than the base. How much malic acid (C4H5O5) to be used? Show calculation; give the result in g

sam (.sam.):

We have acid C4H5O5, and base NaHCO3, It says that we use 1.1 times more acid. we'll use 1.1 later.

OpenStudy (chmvijay):

you have not got his @simoneforver ur trying it from yesterday right

OpenStudy (chmvijay):

sam may help you out

OpenStudy (anonymous):

Yes I hope so :)

OpenStudy (chmvijay):

@.Sam. common answer the @simoneforver question and burn it like anything

sam (.sam.):

So, we can find out the number of moles of acid, n(C4H5O5) using the moles of n(NaHCO3) by just dividing it, According to C4H5O5 + 3NaHCO3 -> C4H2Na3O5+ 3H2O + 3CO2 3 NaHCO3 gives 1 C4H5O5 , so, \[5.94\cancel{molNaHCO_3} \times \frac {1 molC_4H_5O_5}{3\cancel{mol NaHCO_3}}=1.98molC_4H_5O_5\]

OpenStudy (anonymous):

Aa ok then the malic acid in gram is m=1,98g *134g/mol = 265,32g

OpenStudy (anonymous):

No first I have to do this... 1,98mol *1,1 right?

sam (.sam.):

Yes and

OpenStudy (anonymous):

Yes so it is m=(1,98*1,1)mol*134g/mol =

sam (.sam.):

\[2.178mol \times \frac{134g}{1mol}=\]

OpenStudy (anonymous):

Yes :) thanks for your help..

sam (.sam.):

welcome :)